Regina rows upstream at 4mph and then rows back with the current at 6mph. If her time was 2 hours and 15 minutes how far upstream did Regina row?

Answer by Alan3354(69443)

Regina rows upstream at 4mph and then rows back with the current at 6mph. If her time was 2 hours and 15 minutes how far upstream did Regina row?
------------
Avg speed for the round trip = 2*4*6/(4+6) = 4.8 mi/hr
RT distance = 4.8*2.25 = 10.8 miles
One way = 5.4 miles